The measure of the second angle in a triangle is four more than the measure of the first angle and the of the third angle is eight more than twice the measure of the first angle. Find the measure of each angle.

Answer by rfer(16322)

x+x+4+2x+8=180
4x=168
x=42
x+4=46
2x+8=92